OpenStudy (anonymous):

the length of a rectangle is increasing at the rate of 2 m/s, while the width is increasing at a rate of 1 m/s. When the length is 5 m and the width is 3 m how fast is the area increasing?

OpenStudy (bahrom7893):

RELATED RATES

OpenStudy (bahrom7893):

A = l*w dl/dt = 2 dw/dt = 1 l=5 w=3 dA/dt - ?

OpenStudy (bahrom7893):

dA/dt = l*dw/dt + w*dl/dt = 5*1+3*2 = 5+6 = 11

OpenStudy (anonymous):

given dl/dt=2m/s dw/dt=1m/s A=L*W da/dt=ldw/dt+wdl/dt =5*1+3*2=5+6=11 da/dt=11m/s